OpenStudy (anonymous):

Solving systems of linear and quadratic equations using elimination...I can;t figure out how to do these problems,can someone explain please?

OpenStudy (anonymous):

y=-x+3 y=x^2+1

OpenStudy (anonymous):

\[ -x+3=x^2+1 \implies x^2+x-2=0 \]

OpenStudy (anonymous):

OK but how do I get to that answer?That's what I don't understand.

OpenStudy (anonymous):

Equality is transivitve: \[ y=-x+3\\ y=x^2+1\\ -x+3 = y=x^2+ \]

OpenStudy (anonymous):

There are two ways you could do this. Subtract equation 1 from equation 2. (Elimination) Solve for one variable and replace it into the other equation (Substitution)

OpenStudy (anonymous):

Ok,that makes sense,thank you :)

OpenStudy (anonymous):

Ok so if I have the systems,y=x^2 and y+x+2...would they equal x^2-x+2?

OpenStudy (anonymous):

Using substitution on the equations, you get: \[ -x+3=x^2+1 \]Then you subtract \(-x+3\) form both sides:\[ 0 = x^2+1-(-x+3) = x^2+1+x-3=x^2+x-2 \]Or simply put: \[ 0 = x^2+x-2 \]This can be solved with the quadratic formula.

OpenStudy (anonymous):

You don't need the quadratic formula. This can be factored normally.\[x^2+x-2=0 \rightarrow (x+2)(x-1)=0 \rightarrow x=-2,1\] @DogLuverSarah2012